OpenStudy (anonymous):

hey what is 5 over 9 = 3n-2 over 18 i got 10 but the textbook answer says 9 can u please help me

OpenStudy (anonymous):

ok lets see if I can go through and find out where you might have gone wrong

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

ok I got n=4

OpenStudy (anonymous):

But how if the book says 9

OpenStudy (anonymous):

5/9=3n-2/18 multiply each side by 18 10=3n-2 add 2 to each side 12=3n divide each side by 3 4=n

OpenStudy (anonymous):

what do the directions say in the book

OpenStudy (anonymous):

they just say to find what the exponent equals

OpenStudy (anonymous):

well n=4 idk how they got it equals 9..?

OpenStudy (anonymous):

show me what you did to get 10

OpenStudy (anonymous):

3n-2=1n 9*1=9n 9n=5*18 9n=90 90/9=10 n=10

OpenStudy (anonymous):

thats what i did

OpenStudy (anonymous):

ok well first of al you can only add/subtract/divide/multiply common things so you wouldn't be able to subtract 2 from 3n because they are not alike.

OpenStudy (anonymous):

so should i do this: 9-2=7 7*3n=21n 21n=5*18 21n=90 90/27=3 i dont get this

OpenStudy (anonymous):

i think u've read something wrong with the question or the answer, cause with how u've typed it up i agree that the answer would be n=4

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

ok bye